OpenStudy (anonymous):

What is the equation of the line in slope-intercept form? the line parallel to y = –8x – 3 through (–1, 4)

OpenStudy (anonymous):

y = –8x –12 y = x –4 y = 8x –4 y = –8x –4

OpenStudy (mathstudent55):

The slpe-intercept form is y = mx + b where m = slope and b is the y-intercept. Parallel lines have the same slope.

OpenStudy (anonymous):

I dont understand :p

OpenStudy (mathstudent55):

The slope of a line shows its tilt. A positive slope means the line goes up to the right. A negative slope means the line goes down to the right. Parallel lines have the same slope. Since you have a given line with equation y = -8x - 3, the slope is -8. The line you need to find has the same slope so it will have equation y = -8x + b To find b, substitue (-1, 4) in for x and y into y = -8x + b, and solve for b.

OpenStudy (anonymous):

@mathstudent55 i got -4=b

OpenStudy (anonymous):

y=8x-4?

OpenStudy (mathstudent55):

y = –8x + b 4 = -8(-1) + b 4 = 8 + b -4 = b You are correct. Now that you know b = -4, and the slope is -8, use those values in y = mx + b, where m = -8, and b = -4.

OpenStudy (anonymous):

ohso it would be y=-8x-4

OpenStudy (mathstudent55):

Almost. be careful, the slope = m = -8, not 8.

OpenStudy (mathstudent55):

You got it!

OpenStudy (anonymous):

YAY! thanks so much!!!

OpenStudy (mathstudent55):

wlcm